OpenStudy (anonymous):

Can someone help me solve 1/2t-3/4=2t+1/3 ?

OpenStudy (anonymous):

Is that 1/2 t as in "(1/2) times "t" "? Or is it ( 1 / 2t ) ?

OpenStudy (anonymous):

What's the difference? Isn't 1/2 times t and (1/2t) the same thing or am I completely wrong?

OpenStudy (anonymous):

No, they are different. Think of t being 100, as an example. 1/2 times 100 is 50. But 1 / ("2 times 100") is equal to 1 / 200 or 0.005

OpenStudy (anonymous):

Oohh! It's 1/2 times t.

OpenStudy (anonymous):

Sorry, I stepped away.

OpenStudy (anonymous):

(1/2)t - 3/4 = 2t + (1/3) (1/2)t -3/4 - (1/3) = 2t (subtract (1/3) term from both sides to get all #s on left) -3/4 - 1/3 = 2t - (1/2)t (subtract (1/2)t from both sides to get all "t" terms on right) -9/12 - 4/12 = (4/2)t - (1/2)t common denominators -13/12 = (3/2)t simplify terms and #s (2/3) (-13/12) = (2/3)(3/2)t multiply both sides by (2/3) to eliminate fraction in front of the "t" -26/36 = t multiply out the left side t = -26/36 = -13/18 reduce the fraction.

OpenStudy (anonymous):

WOW! thanks! that's alot!